Find the probability and interpret the results. If​ convenient, use technology to find the probability. The population mean annual salary for environmental compliance specialists is about $65,000. A random sample of 32 specialists is drawn from this population. What is the probability that the mean salary of the sample is less than $62,000? Assume sigma= ​$5,500.

The probability that the mean salary of the sample is less than $62,000 is ____
Answer #1

Solution :

Given that ,

mean = \mu = 65000

standard deviation = \sigma = 5500

n = 32

\mu\bar x = 65000

\sigma\bar x =\sigma  / \sqrt n = 5500 / \sqrt 32=972.3

P(\bar x < 62000) = P[(\bar x - \mu \bar x ) / \sigma \bar x < (62000-65000) / 972.3]

= P(z <-3.09 )

Using z table

probability= 0.0010